OpenStudy (anonymous):

If f(x) = 2x + 6, find f(7). 14 15 20 6

OpenStudy (anonymous):

2*7+6=20

OpenStudy (anonymous):

Put x = 7 in the function: f(7) = 2*7 + 6 f(7) = 14 + 6 Add and get the answer..